While helping a friend shift furniture you do 2.2 MJ of mechanical work at and efficiency of 11% (η = 0.11).
How much food energy would you need to ingest in order to prevent weight loss? (to 2 s.f and in MJ)
1
Expert's answer
2020-05-21T13:47:19-0400

Solution:

"A=2.2 MJ;"

"\eta=0.11;"

"Q-?;"

The efficiency is calculated by the formula:

"\eta=\dfrac{A}{Q};\implies Q=\dfrac{A}{\eta};"

"Q=\dfrac{2.2MJ}{0.11}=20MJ;"

Answer: "Q=20MJ."